o f juniors and seni Of the 5 out if they pro Total Hamburger Hot Dog 13 53 40 A group of students, consisting of juniors and both if possible! thank you!
Answer #1

question 1

a) total students (N) = 53+37 = 90

b) students who preferred hot dogs = 13 + 13 = 26

c) N(Junior n Hot dog) = 13

d) N(Senior n hamburger) = 24

e) N(Junior U Hamburger) = N(Junior) + N(hamburger) - N(Junior n Hamburger) = 53 + 64 - 40 = 77

f) N(Senior n Hot dog) = 13

g) N(not hamburger) = N - N(student eating hamburger) = 90 - 64 = 26

N(not Junior) = N - N(Junior) = 90 - 53 = 37